OpenStudy (anonymous):

Anthony is making a collage for his art class by picking shapes randomly. He has five squares, two triangles, two ovals, and four circles. Find P(a triangle or a square is chosen first).

OpenStudy (kropot72):

There total quantity of shapes is 5 + 2 + 2 + 4 = 13. The total quantity of triangles and squares is 2 + 5 = 7. The probability of choosing a triangle or a square first is given by \[P(\triangle\ or\ \square\ chosen\ first)=\frac{7}{13}\]